# Functional Natural Policy Gradients

### 30 Mar 2026 v1

@@ 1 Introduction
A variety of work studies semiparametric efficient estimation of the value of a fixed policy from offline data (Scharfstein et al., 1999; Chernozhukov et al., 2018; Dudík et al., 2011; Jiang and Li, 2016; Kallus and Uehara, 2020, 2022; Kallus et al., 2022). And, a variety of work considers selecting the policy that optimizes such estimates over policies in a given class (Athey and Wager, 2021; Zhang et al., 2013; Kallus, 2021; Zhou et al., 2023; Foster and Syrgkanis, 2023; chernozhukov2019semi), which generally yields rates the scale with policy class complexity, e.g., $O_{P}(N^{-1/2})$ for VC classes. Luedtke and Chambaz (2020) get regret acceleration to $o_{P}(N^{-1/2})$ by leveraging an equicontinuity argument. Margin (or, noise) conditions controlling the fraction of hard decision contexts can also speed up policy-search regret (Boucheron et al., 2005; Hu et al., 2022, 2024; Tsybakov, 2004; Tsybakov and van de Geer, 2005; Koltchinskii, 2006). However, with the exception of Bennett and Kallus (2020) whose assumptions imply the optimal policy solves a smooth conditional moment restriction, none of these works leverage cross-fitting or debiasing to estimate the optimal policy itself, such as done for example when targeting a low-dimensional regular functional (Chernozhukov et al., 2024).
@@
We demonstrate how to solve for the functional value-optimality condition via optimization along a one-dimensional policy subclass, which we construct in the same way van2016one construct a universal-least favorable submodel (ULFM) within a nuisance space. It turns out that, in the policy optimization setting, ULFMs are natural policy gradient (NPG) flows (Kakade, 2001). In that sense, our proposal can be seen as a functional NPG ascent, and is similar in spirit to functional optimization (see Petrulionyte et al., 2024; khoury2025learning for example).
@@ 3 Semiparametric natural gradient and natural gradient flow
Our construction mimics the least favorable uniform submodel construction in van2016one.

### 03 Apr 2026 v2

@@ 1 Introduction
A variety of work studies semiparametric efficient estimation of the value of a fixed policy from offline data (Scharfstein et al., 1999; Chernozhukov et al., 2018; Dudík et al., 2011; Jiang and Li, 2016; Kallus and Uehara, 2020, 2022; Kallus et al., 2022). And, a variety of work considers selecting the policy that optimizes such estimates over policies in a given class (Athey and Wager, 2021; Zhang et al., 2013; Kallus, 2021; Zhou et al., 2023; Foster and Syrgkanis, 2023; Chernozhukov et al., 2019), which generally yields rates the scale with policy class complexity, e.g., $O_{P}(N^{-1/2})$ for VC classes. Luedtke and Chambaz (2020) get regret acceleration to $o_{P}(N^{-1/2})$ by leveraging an equicontinuity argument. Margin (or, noise) conditions controlling the fraction of hard decision contexts can also speed up policy-search regret (Boucheron et al., 2005; Hu et al., 2022, 2024; Tsybakov, 2004; Tsybakov and van de Geer, 2005; Koltchinskii, 2006). However, with the exception of Bennett and Kallus (2020) whose assumptions imply the optimal policy solves a smooth conditional moment restriction, none of these works leverage cross-fitting or debiasing to estimate the optimal policy itself, such as done for example when targeting a low-dimensional regular functional (Chernozhukov et al., 2024).
@@
We demonstrate how to solve for the functional value-optimality condition via optimization along a one-dimensional policy subclass, which we construct in the same way van der Laan and Gruber (2016) construct a universal-least favorable submodel (ULFM) within a nuisance space. It turns out that, in the policy optimization setting, ULFMs are natural policy gradient (NPG) flows (Kakade, 2001). In that sense, our proposal can be seen as a functional NPG ascent, and is similar in spirit to functional optimization (see Petrulionyte et al., 2024; El Khoury et al., 2025 for example).
@@ 3 Semiparametric natural gradient and natural gradient flow
Our construction mimics the least favorable uniform submodel construction in van der Laan and Gruber (2016).